Laser Absorption Spectroscopy of LaF^{+} : Accurate Term Energies for the H^{2}Π, D^{2}Φ, and X^{2}Δ States

L. A. Kaledin,\underline{A. L. Kaledin}, and M. C. Heaven, Department of Chemistry, Emory University, Atlanta, GA 30322.

The laser absorption spectrum of the LaF^{+} molecular ion has been observed for the first time. Gas phase LaF was obtained by heating a mixture of a 0.2 g of La metal and 0.3 g of LaF_{3} powder to 1900 K. LaF^{+} ions were produced by passing a 25 kHz discharge (80-100W, 0.5A peak current) through the vapor. Under these conditions the Tantalum heating tube used to vaporize the sample also acted as a hollow cathode. 138 lines of the ^{2}Φ_{2.5}-^{2}Δ_{1.5} (0,0) band were recorded. This band system was observed previously ^{a} in the emission spectrum of LaF^{+}. Accurate term energies are (in cm^{−1}):

\begin{matrix}T_{0}(H^{2}Π_{1.5}) = & 30398.78(5), & T_{0}(D^{2}Φ_{3.5}) = & 17099.24(5), & T_{0}(X^{2}Δ_{2.5}) = & 274.62(5), \\ T_{0}(H^{2}Π_{0.5}) = & 30237.85(5), & T_{0}(D^{2}Φ_{2.5}) = & 16534.738(5), & T_{0}(X^{2}Δ_{1.5}) = & 0.0 \end{matrix}

Ligand field theory models were used to provide the configurational assignment for the LaF^{+} electronic states. The symmetry for the ground state of the LaF^{+} molecular ion has been reassigned as X^{2}Δ.

Work supported by AFOSR under grant F19628-90-C-025.

aE. A. Shenyavskaya and L. V. Gurvich, J. Mol. Spectrosc. 81, 152-163 (1980).